# Neural Networks and Friction: Slide, Hold, Learn

### 21 Feb 2024 v1

@@ III Training results
The overall training procedure is depicted in fig. 7 and the final losses are consigned in table 1. The complex loss function eq. 13 was not used at once, but an incremental approach was explored: first, the simplest loss (mean squared error, $\mathsf{ME}$) was used to initialize weights and biases. Second, the network was retrained adding the initial-condition term ($\mathsf{ME+ics}$), which further reduced the loss by 50%. This main part was again extracted and used into a larger scheme that included the direct effect loss term ($\mathsf{ME+ics+de}$). The overall precision suffered, but, as it will be shown, this deterioration of the objective function minimization does not translate into a substantial deterioration of the overall approximation; rather, it seems that this larger value reflect the struggle to capture the jumps’ magnitude. We tried re-training the network from $\mathsf{ME+ics}$ for both direct effect and healing simultaneously ($\mathsf{ME+ics+de+heal}$), to no avail. Apparently, the addition of the healing phenomenology to the loss further deflect from the optimum, but it is believed that this is really a drawback of the RSF model used to generate the synthetic data.
Figure 7: Mean training loss: $\mathsf{ME}$ is trained without initialization, $\mathsf{ME+ics}$ is trained with paramters from $\mathsf{ME}$, $\mathsf{ME+ics+de}$ and $\mathsf{ME+ics+de+heal}$ are trained with parameters from $\mathsf{ME+ics}$.
Table 1: Final loss value for different loss functions: ME refers to mean squared error between predictions and targets, ICs refers to enforcing the first value of predictions to be equal to targets’, DE does to enforces friction jumps in target data and predictions to be of the same magnitude, H refers to enforcing logarithmic healing.
@@ IV Test
Once the training is complete, we take the main net, which is our ready-to-use $\mathsf{NN}$ function that is to model the interface evolution. We test it using the 3000 sequences that the net did not see during training. For each sequence, the percent error is computed as
$\mathsf{error}(\%)=100{||\mathsf{NN}(\bm{V}/V_{ch})-\delta\bm{\hat{\mu}}/\mu_{ch}||_{2}\over||\delta\bm{\hat{\mu}}/\mu_{ch}||_{2}}\,.$ (15)
@@ V Discussion
As hinted by the test results presented in table 2, figs. 10 and 10 clearly showcases that the GRU-based network $\mathsf{NN}$ is able of closely reproducing the friction dynamics encapsulated in the datasets generated with a customary rate-and-state phenomenological law.
@@
The fact that, out of the hold periods, the target (friction coefficient) changes when there is no change in the features (loading velocities) is a clear drawback, no matter how slightly the change may be.
@@
which can interfere with the scaling, unless $\theta(t=0)\ll t$. The latter condition may be met in real experiments, in which the holds tend to last for much longer than the characteristic age of the interface, but it is certainly not the case here as the holds are always “short”, i.e., of duration similar magnitude to the magnitude of the state variable. This seems to be the most obvious drawback of not considering explicitly the interface state in the data-driven framework. In any case, it does not appear as insurmountable, since (as justified above)
Yet another aspect to acknowledge is that the magnitude of direct effect peaks seems to be affected by the holds. See that in fig. 10 there is a relatively short hold period preceded and ended by jumps which are well captured by the nets; conversely, the jumps’ magnitude are significantly underpredicted in fig. 10, which features a longer hold, which particularly the net that accounts explicitly for healing struggles to capture.
@@ VI Final Remarks
However, our investigation also identified limitations, particularly concerning the handling of logarithmic healing during short hold periods. This limitation suggests that future models may incorporate mechanisms to better account for short holds, thereby enhancing their accuracy and applicability to a broader range of slip events. Additionally, we observed a potential conflict between optimizing for healing and for the direct effect within the training process, indicating an area for further refinement in how these fundamental aspects of frictional behavior are integrated into the learning scheme. And let alone that no noise has been included in the training datasets yet, and only protocols made of sudden velocity jumps changes were considered.
@@
This document contains but an exploratory approach to modeling frictional interface via deep learning. Looking forward, there are several promising avenues for enhancing the predictive power and generalizability of recurrent neural networks for dynamic friction. Improvement could be sought through the exploration of alternative RNN architectures and more nuanced physics-informed loss functions (for instance, to account for friction manifesting indirectly in the dynamics of spring-block systems [2]). Further parametric studies in terms of network hyperparameters are also warranted, in particular in terms of volumes of available data: unlike cheap synthetic data, experimental data remains relatively scarce; hence it is worth exploring how new techniques [48, 49], aimed at improving data economy in the context of constitutive modeling with neural networks,
@@
As mentioned above, incorporating spring-block dynamics into the model could pave the way for learning from real experimental data, thus circumventing the gap between theoretical models and practical observations. Combined with the use of automatic differentiation [47] (not explored at this point), could offer a completely new pathway to take advantage of experimental data to understand friction, replacing the use of experiments to simply fit parameters by producing training data that is the infused in a network that can be used later for predictions.
@@
Moreover, the exploration of advanced machine learning features such as transformers [50] and Long Short-Term Memory (LSTM) networks [51] could offer additional improvements. These technologies, known for their ability to capture long-distance dependencies within data, might provide the key to overcoming the current limitations. The promise of deep learning does not stop there: one can foresee that an approach based on neural operators [52] may help also with the problem of generalizing friction from lab conditions to crustal ones, the network operator may be brought to bear in this case.
The ultimate objective is to transition from conventional phenomenological models of dynamic friction to more faithful representations directly learned from the same data traditionally used for parameter fitting. This shift should not entail a departure from our pursuit of a fundamental understanding of friction dynamics. On the contrary, it necessitates a more integrated approach where firm physical principles are embedded within the deep learning framework. The adoption of these advanced models should be driven by their proven efficacy in capturing interface behaviors, thereby enhancing their utility in larger-scale simulations where interface response plays a role. This strategy aims not to replace the foundational theories of friction but to complement them with data-driven insights, ensuring a more comprehensive and precise depiction of frictional phenomena to be included in numerical analyses.
@@ Acknowledgements
The author acknowledges financial support from the Swiss National Science Foundation, via Ambizione Grant PZ00P2_216341 “Data-Driven Computational Friction”. Useful conversations (regarding dynamic friction, rate-and-state friction and/or laboratory experiments) with Dr. Gabriel Meyer, Dr. Simon Guérin-Marthe and Ms. Roxane Ferry are gratefully acknowledged. An EPFL bachelor project conducted by Mr. Antoine Binggeli provided Tensorflow code that inspired the use of GRUs to this particular problem.

### 23 Feb 2024 v2

@@ III Training results
The overall training procedure is depicted in fig. 7 and the final losses are consigned in table 1. The complex loss function eq. 13 was not used at once, but an incremental approach was explored: first, the simplest loss (mean squared error, $\mathsf{ME}$) was used to initialize weights and biases. Second, the network was retrained adding the initial-condition term ($\mathsf{ME+ics}$), which further reduced the loss by 50%. This main part was again extracted and used into a larger scheme that included the direct effect loss term ($\mathsf{ME+ics+de}$). The overall precision suffered, but, as it will be shown, this deterioration of the objective function minimization does not translate into a substantial deterioration of the overall approximation; rather, it seems that this larger value reflects the struggle to capture the jumps’ magnitude. We tried re-training the network from $\mathsf{ME+ics}$ for both direct effect and healing simultaneously ($\mathsf{ME+ics+de+heal}$), to no avail. Apparently, the addition of the healing phenomenology to the loss further deflects from the optimum, but it is believed that this is really a drawback of the RSF model used to generate the synthetic data.
Figure 7: Mean training loss: $\mathsf{ME}$ is trained without initialization, $\mathsf{ME+ics}$ is trained with parameters from $\mathsf{ME}$, $\mathsf{ME+ics+de}$ and $\mathsf{ME+ics+de+heal}$ are trained with parameters from $\mathsf{ME+ics}$.
Table 1: Final loss value for different loss functions: ME refers to mean squared error between predictions and targets, ICs refers to enforcing the first value of predictions to be equal to targets’, DE does to enforcing friction jumps in target data and predictions to be of the same magnitude, H refers to enforcing logarithmic healing.
@@ IV Test
Once the training is complete, we take out the main net, which is our ready-to-use $\mathsf{NN}$ function that is to model the interface evolution. We test it using the 3000 sequences that the net did not see during training. For each sequence, the percent error is computed as
$\mathsf{error}(\%)=100{||\mathsf{NN}(\bm{V}/V_{ch})-\delta\bm{\hat{\mu}}/\delta\mu_{ch}||_{2}\over||\delta\bm{\hat{\mu}}/\delta\mu_{ch}||_{2}}\,.$ (15)
@@ V Discussion
As hinted by the test results presented in table 2, figs. 10 and 10 clearly showcase that the GRU-based network $\mathsf{NN}$ is able of closely reproducing the friction dynamics encapsulated in the datasets generated with a customary rate-and-state phenomenological law.
@@
The fact that the target (friction coefficient) changes during the steady state when there is no change in the features (loading velocities) is a clear drawback, no matter how slightly the change may be.
@@
which can interfere with the scaling, unless $\theta(t=0)\ll t$. The latter condition may be met in real experiments, in which the holds tend to last for much longer than the characteristic age of the interface, but it is certainly not the case here as the holds are always “short”, i.e., of duration similar to the magnitude of the state variable. This seems to be the most obvious drawback of not considering explicitly the interface state in the data-driven framework. In any case, it does not appear as insurmountable, since (as justified above) in real data the behavior during periods of rest would adjust better to the logarithmic scaling.
Yet another aspect to acknowledge is that the magnitude of direct effect peaks seems to be affected by the holds. See for instance that in fig. 10 there is a relatively short hold period preceded and ended by jumps which are well captured by the nets; conversely, the jumps’ magnitude are significantly underpredicted in fig. 10, which features a longer hold, particularly by the net tuned explicitly for healing.
@@ VI Final Remarks
However, our investigation also identified limitations, particularly concerning the handling of logarithmic healing during hold periods. Additionally, we observed a potential conflict between optimizing for healing and for the direct effect within the training process, indicating an area for further refinement in how these fundamental aspects of frictional behavior are integrated into the learning scheme. And let alone that no noise has been included in the training datasets yet, and only protocols made of sudden velocity jumps changes were considered.
@@
This document contains but an exploratory approach to modeling frictional interface via deep learning. Looking forward, there are several promising avenues for enhancing the predictive power and generalizability of recurrent neural networks for dynamic friction. Improvement could be sought through the exploration of alternative RNN architectures and more nuanced physics-informed loss functions (for instance, to account for friction manifesting indirectly in the dynamics of spring-block systems [2]). Further parametric studies in terms of network hyperparameters are also warranted, in particular in terms of volumes of available data: unlike cheap synthetic data, experimental data remains relatively scarce; hence it is worth exploring how new techniques [54, 55], aimed at improving data economy in the context of constitutive modeling with neural networks, can help in this case.
@@
As mentioned above, incorporating spring-block dynamics into the model could pave the way for learning from real experimental data, thus circumventing the gap between theoretical models and practical observations. Combined with the use of automatic differentiation [53] (not explored at this point), it could offer a completely new pathway to take advantage of laboratory restuls to understand friction, replacing the use of experiments to simply fit parameters by producing training data that is the infused in a network that can be used later for predictions.
@@
Moreover, the exploration of advanced machine learning features such as transformers [56] and Long Short-Term Memory (LSTM) networks [57] could offer additional improvements. These technologies, known for their ability to capture long-distance dependencies within data, might provide the key to overcoming some of the current limitations. The promise of deep learning does not stop there: one can foresee that an approach based on neural operators [58] may help also with the problem of generalizing friction from lab conditions to crustal ones.
The end goal is to evolve from traditional phenomenological friction models towards more accurate representations that are likewise derived from the data already used for parameter fitting. This evolution does not entail giving up on a deeper understanding of friction dynamics. Rather, it calls for a complementary integration of solid physical principles with the deep learning paradigm. The rationale for adopting these deep-learning models shall be their demonstrable ability to accurately capture complex interface physics, which, in turn, would justify their use in larger-scale numerical simulations involving those phenomena.
@@ Acknowledgements
The author acknowledges financial support from the Swiss National Science Foundation, via Ambizione Grant PZ00P2_216341 “Data-Driven Computational Friction”. Useful conversations (regarding text improvements, rate-and-state friction and/or laboratory experiments) with Dr. Gabriel Meyer, Dr. Simon Guérin-Marthe and Ms. Roxane Ferry are gratefully acknowledged. An EPFL bachelor project conducted by Mr. Antoine Binggeli provided Tensorflow code that inspired the use of GRUs to this particular problem.